# Multiperiodic Processes: Ergodic Sources with a Sublinear Entropy

### v2

@@ 2.2 Multiperiodic processes
Definition 2 (multiperiodic process)
@@
  • Proof: Observe that events $(\Sigma_{k}=1)$ are independent. Hence by the Borel-Cantelli lemmas [24, Theorems 4.3 and 4.4], set $\left\{k\in\mathbb{N}:\Sigma_{k}=1\right\}$ is infinite almost surely if and only if
@@
    In view of the uniform distributions of $(\Sigma_{k})_{k\in\mathbb{N}}$, we obtain the claim. $\Box$
@@
  • Proof: A sequence $(a_{t})_{t\in\mathbb{Z}}$ is called $p$-periodic if
@@
    for a certain $p$-periodic sequence $(a_{t})_{t\in\mathbb{Z}}$. It is known that $p$-periodic processes are stationary and ergodic but they are not mixing [5, Section 4.1].
@@
    Thus process $(K^{\leq m}_{t})_{t\in\mathbb{Z}}$ is stationary and ergodic but is not mixing. Since $m\in\mathbb{N}$ is arbitrarily chosen, hence also process $(K_{t})_{t\in\mathbb{Z}}$ is stationary and ergodic but is not mixing. The reason is that verifying stationarity, ergodicity, and mixing can be reduced to checking probabilities of finite blocks by [5, Theorem 4.10] or [25, Lemma 7.15]. $\Box$
@@ 3.1 Relative frequencies
$\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{t}{\bf 1}{\left\{k_{j}\geq k\right\}}=f_{k}:=\prod_{j=1}^{k-1}\frac{\pi_{j}-1}{\pi_{j}}.$ (21)
  • Proof: Let $(k_{t}^{\geq r})_{t\in\mathbb{N}}$ be the subsequence of $(k_{t})_{t\in\mathbb{N}}$ from which we have removed all tokens $k_{t}<r$. We proceed by induction on $k$. First, we observe
    $\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{t}{\bf 1}{\left\{k_{j}\geq 1\right\}}=1=f_{1}.$ (22)
@@
    $\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{t}{\bf 1}{\left\{k_{j}\geq k\right\}}$
    $=\left(\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{t}{\bf 1}{\left\{k_{j}\geq k-1\right\}}\right)\left(\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{t}{\bf 1}{\left\{k_{j}^{(k-1)}\geq k\right\}}\right)$
@@
    Hence the claim follows. $\Box$
Theorem 6
@@
  • Proof: We obtain distribution (24) by the Birkhoff ergodic theorem, which establishes the almost sure equality of probabilities and relative frequencies for stationary ergodic processes [26, 27]. $\Box$
@@ 3.2 Waiting times
  • Proof: Consider a prefix $(K_{1},K_{2},\ldots,K_{T_{1}})$ of the multiperiodic process. When we delete all $K_{t}<r$, the sequence shortens to $(K^{\geq r}_{1},K^{\geq r}_{2},\ldots,K^{\geq r}_{T_{r}})$. We observe that $W_{k}\leq T_{1}$ if $T_{k}=\pi_{k}$, whereas the partial periodicity implies
@@
Definition 3 (increasing periods)
@@
  • Proof: We have $w_{k+1}\geq w_{k}$ since
@@ 3.3 Number of types
$V_{t}:=\operatorname{\#}\mathcal{V}_{t}.$ (37)
Let us introduce also three more statistics:
@@
  • Proof: The inequalities follow by $w_{k}\geq W_{k}$ and $U_{t}-1\leq\operatorname{\#}\mathcal{V}_{t}$. $\Box$
Applying the idea by [14] for memoryless sources, we can prove this proposition.
Theorem 10
@@
  • Proof: We observe
    $\operatorname{\mathbf{E}}V_{t}=\operatorname{\mathbf{E}}\sum_{k=1}^{\infty}{\bf 1}{\left\{k\in\mathcal{V}_{t}\right\}}=\sum_{k=1}^{\infty}P(k\in\mathcal{V}_{t})\leq g_{t}+\sum_{k>g_{t}}P(k\in\mathcal{V}_{t})$ (44)
@@
    Hence the claim follows. $\Box$
@@
  • Proof: We observe that bound
@@
    holds for any function $h_{t}$. In particular, for an $\epsilon>0$, we may take $h_{t}=\epsilon t/2$. For all sufficiently large $t$, we observe $P(K_{0}>h_{t})\leq\epsilon/2$. Hence, for these $t$, we have $\operatorname{\mathbf{E}}V_{t}/t\leq h_{t}/t+P(K_{0}>h_{t})\leq\epsilon$. By arbitrariness of $\epsilon$, we derive (46). $\Box$
@@
[…]
  • Proof: Observe by the Markov inequality that
@@
    Hence $\lim_{k\to\infty}\operatorname{\mathbf{E}}V_{2^{k}}k/2^{k}=0$, whereas by the Borel-Cantelli lemma [24, Theorems 4.3], we derive $V_{2^{k}}k/2^{k}<\epsilon$ for all but finitely many $k$ almost surely. Notice now that $V_{t}$ is a growing function of $t$. Hence $(V_{t}\log t)/t\leq 2V_{2^{k}}k/2^{k}$ for $2^{k-1}\leq t\leq 2^{k}$. In particular, we have (48) and $\limsup_{t\to\infty}(V_{t}\log t)/t\leq 2\epsilon$ almost surely. By arbitrariness of $\epsilon$, we derive (49). $\Box$
@@ 4.1 Seed estimator
  • Proof:
@@
    • If $k\in\mathcal{V}_{t}$ then $k\leq\max\mathcal{V}_{t}$ and $\Sigma_{k}=L_{k}\leq t$ since the first $L_{k}$ tokens of sequence $(K^{(k)}_{t})_{t\in\mathbb{N}}$ are a subsequence of sample $K_{1}^{t}$. Hence we obtain claim 4.
    $\Box$
@@ 4.2 Block entropy
  • Proof: Let $p>1$ and $q=p/(p-1)$. For $X,Z\geq 0$, the Hölder inequality is
@@
    Using $Z=\ln Y$ in the Hölder inequality and chaining it with the above chain of inequalities, we derive the claim. $\Box$
@@
  • Proof: By claim 1. of Theorem 13 and property $k<u_{t}\implies k\in\mathcal{V}_{t}$, the tuple of seeds $\Sigma_{1}^{u_{t}-1}$ is a function of sample $K_{1}^{t}$. Hence we obtain the lower bound
@@
    which is the claimed inequality. $\Box$
@@
  • Proof: The claim follows by (48) and (59) since $\operatorname{\mathbf{E}}K_{0}^{1/q}<\infty$ for a $q\geq 1$ implies $\operatorname{\mathbf{E}}(\log K_{0})^{3}<\infty$. $\Box$
@@ 5.2 Linear periods
[…]
@@ 6 Conclusion
Our toy example of multiperiodic processes extends mathematical analysis of Hilberg‘s law—the power law growth of redundancy for certain stationary sources, possibly entailing statistical language models [2, 3]. This topic is relevant not only for quantitative linguistics [4, 5] but also for large language models in machine learning, via the neural scaling law [6, 7, 8]. As it was shown in the previous research revolving around the example of Santa Fe processes [4, 5, 18, 10], the subject can be linked to traditional analyses of Zipf‘s law [11, 12]. This study has supplied further findings in this vein.
@@ Acknowledgment
I thank Jan Mielniczuk, Szymon Jaroszewicz, Marcus Hutter, and Guido Montúfar for discussing prior versions of the paper.

### v3

@@ 2.2 Multiperiodic processes
Definition 2 (multiperiodic process).
@@
Proof.
Observe that events $(\Sigma_{k}=1)$ are independent. Hence by the Borel-Cantelli lemmas [5, Theorems 4.3 and 4.4], set $\left\{k\in\mathbb{N}:\Sigma_{k}=1\right\}$ is infinite almost surely if and only if
@@
In view of the uniform distributions of $(\Sigma_{k})_{k\in\mathbb{N}}$, we obtain the claim. ∎
@@
Proof.
A sequence $(a_{t})_{t\in\mathbb{Z}}$ is called $p$-periodic if
@@
for a certain $p$-periodic sequence $(a_{t})_{t\in\mathbb{Z}}$. It is known that $p$-periodic processes are stationary and ergodic but they are not mixing [24, Section 4.1].
@@
Thus process $(K^{\leq m}_{t})_{t\in\mathbb{Z}}$ is stationary and ergodic but is not mixing. Since $m\in\mathbb{N}$ is arbitrarily chosen, hence also process $(K_{t})_{t\in\mathbb{Z}}$ is stationary and ergodic but is not mixing. The reason is that verifying stationarity, ergodicity, and mixing can be reduced to checking probabilities of finite blocks by [24, Theorem 4.10] or [37, Lemma 7.15]. ∎
@@ 3.1 Relative frequencies
$\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{t}{\mathbf{1}}{\left\{k_{j}\geq k\right\}}=f_{k}:=\prod_{j=1}^{k-1}\frac{\pi_{j}-1}{\pi_{j}}.$ (25)
Proof.
Let $(k_{t}^{\geq r})_{t\in\mathbb{N}}$ be the subsequence of $(k_{t})_{t\in\mathbb{N}}$ from which we have removed all tokens $k_{t}<r$. We proceed by induction on $k$. First, we observe
$\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{t}{\mathbf{1}}{\left\{k_{j}\geq 1\right\}}=1=f_{1}.$ (26)
@@
$\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{t}{\mathbf{1}}{\left\{k_{j}\geq k\right\}}$
$=\left(\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{t}{\mathbf{1}}{\left\{k_{j}\geq k-1\right\}}\right)\left(\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{t}{\mathbf{1}}{\left\{k_{j}^{(k-1)}\geq k\right\}}\right)$
@@
Hence the claim follows. ∎
Theorem 6.
@@
Proof.
We obtain distribution (28) by the Birkhoff ergodic theorem, which establishes the almost sure equality of probabilities and relative frequencies for stationary ergodic processes [6, 34]. ∎
@@ 3.2 Waiting times
Proof.
Consider a prefix $(K_{1},K_{2},\ldots,K_{T_{1}})$ of the multiperiodic process. When we delete all $K_{t}<r$, the sequence shortens to $(K^{\geq r}_{1},K^{\geq r}_{2},\ldots,K^{\geq r}_{T_{r}})$. We observe that $W_{k}\leq T_{1}$ if $T_{k}=\pi_{k}$, whereas the partial periodicity implies
@@
Definition 3 (increasing periods).
@@
Proof.
We have $w_{k+1}\geq w_{k}$ since
@@ 3.3 Number of types
$V_{t}:=\operatorname{card}\mathcal{V}_{t}.$ (41)
Let us introduce also four more statistics:
@@
Proof.
The inequalities follow by $w_{k}\geq W_{k}$ and $U_{t}-1\leq\operatorname{card}\mathcal{V}_{t}$. ∎
Generalizing an idea used by Khmaladze [48] for memoryless sources, we can prove this proposition.
Theorem 10.
@@
Proof.
We observe
$\operatorname{\mathbf{E}}V_{t}=\operatorname{\mathbf{E}}\sum_{k=1}^{\infty}{\mathbf{1}}{\left\{k\in\mathcal{V}_{t}\right\}}=\sum_{k=1}^{\infty}P(k\in\mathcal{V}_{t})\leq g_{t}+\sum_{k>g_{t}}P(k\in\mathcal{V}_{t})$ (48)
@@
Hence the claim follows. ∎
@@
Proof.
Applying the ideas from the proof of the previous theorem, we observe that bound
@@
holds for any function $h_{t}$. In particular, for an $\epsilon>0$, we may take $h_{t}=\epsilon t/2$. For all sufficiently large $t$, we observe $P(K_{0}>h_{t})\leq\epsilon/2$. Hence, for these $t$, we have $\operatorname{\mathbf{E}}V_{t}/t\leq h_{t}/t+P(K_{0}>h_{t})\leq\epsilon$. By arbitrariness of $\epsilon$, we derive (50). ∎
@@
$\lim_{t\to\infty}\frac{V_{t}\log t}{t}$ $=0\text{ a.s.}$ (53)
Proof.
Observe by the Markov inequality that
@@
Hence $\lim_{k\to\infty}\operatorname{\mathbf{E}}V_{2^{k}}k/2^{k}=0$, whereas by the Borel-Cantelli lemma [5, Theorems 4.3], we derive $V_{2^{k}}k/2^{k}<\epsilon$ for all but finitely many $k$ almost surely. Notice now that $V_{t}$ is a growing function of $t$. Hence $(V_{t}\log t)/t\leq 2V_{2^{k}}k/2^{k}$ for $2^{k-1}\leq t\leq 2^{k}$. In particular, we have (52) and $\limsup_{t\to\infty}(V_{t}\log t)/t\leq 2\epsilon$ almost surely. By arbitrariness of $\epsilon$, we derive (53). ∎
@@ 4.1 Seed estimator
Proof.
The subsequent claims are dealt with below.
@@
  • If $k\in\mathcal{V}_{t}$ then $k\leq\max\mathcal{V}_{t}$ and $\Sigma_{k}=L_{k}\leq t$ since the first $L_{k}$ tokens of sequence $(K^{(k)}_{t})_{t\in\mathbb{N}}$ are a subsequence of sample $K_{1}^{t}$. Hence we obtain claim 4. ∎
@@ 4.2 Block entropy
Proof.
Let $p>1$ and q=p/(p1)q=p/(p-1). For $X,Z\geq 0$, the Hölder inequality is
@@
Using $Z=\ln Y$ in the Hölder inequality and chaining it with the above chain of inequalities, we derive the claim. ∎
@@
Proof.
By claim 1. of Theorem 13 and property $k<u_{t}\implies k\in\mathcal{V}_{t}$, the tuple of seeds $\Sigma_{1}^{u_{t}-1}$ is a function of sample $K_{1}^{t}$. Hence we obtain the lower bound
@@
which is the claimed inequality. ∎
@@
Proof.
The claim follows by (52) and (63) since $\operatorname{\mathbf{E}}K_{0}^{1/q}<\infty$ for a $q\geq 1$ implies $\operatorname{\mathbf{E}}(\log K_{0})^{3}<\infty$. ∎
@@ 5.2 Linear periods
$\operatorname{\mathbf{E}}K_{t}^{1/q}$ $=\int_{0}^{\infty}P(K_{t}^{1/q}\geq k)dk=\int_{0}^{\infty}P(K_{t}\geq k^{q})dk$
$=\int_{0}^{\infty}f_{\left\lceil k^{q}\right\rceil}dk\leq\int_{1}^{\infty}\left(\frac{c+2}{c}\right)^{1/c}k^{-q/c}dk<\infty.$ (87)
@@ 6 Conclusion
The toy example of multiperiodic processes extends mathematical analysis of Hilberg‘s law—the power law growth of redundancy for certain stationary sources, possibly entailing statistical language models. This topic is relevant not only for quantitative linguistics but also for large language models in machine learning, via the neural scaling law. As it was shown in the previous research revolving around the example of Santa Fe processes, the topic can be linked to more traditional analyses of Zipf‘s law. This study has supplied further findings in this vein.
@@ Acknowledgment
I thank Jan Mielniczuk, Szymon Jaroszewicz, Marcus Hutter, and Guido Montúfar for discussing prior versions of this paper.